Question 4(4 points Two vectors have the following components A (0.4,3,7) and B8.6,3.9,What is th magnitude of the sum of the two vectors? Your Answer Answer Question 5(3 points The X component of a vectoris 5.6 and the Y component is 10.0.What s he angle between he positiveXaxis and the vector Your Answer Answer
Answer #1

Solution 4)

Given, A(0.4,3.7) and B(8.6,3.9)

Now, resultant R= (0.4 + 8.6) i + (3.7+3.9)j = 9i + 7.6 j

Magnitude= sqrt( 9^2 + 7.6^2) = 11.77 (Ans)

=======

Solution 5)

Theta= tan^-1(10/5.6) = 60.75 degrees (Ans)

=======